25.0 cm3 of 0.100 moldm-3 sodium carbonate solution was titrated with dilute phosphoric acid of unknown concentration. 17.60 cm3 of the acid was required to neutralise the sodium carbonate. Find the concentration of the phosphoric acid in moldm-3 and gdm-3
1
Expert's answer
2020-12-21T03:38:16-0500

3Na2CO3 + 2H3PO4 = 2Na3PO4 + 3H2O + 3CO2

Ca Va / CbVb = na / nb

Ca × 17.60 /(0.1 ×25) = 2/3

Ca = 0.1 × 25 ×2 /(17.60×3)

= 0.095M

Mass concentration = molar concentration × molar mass

= 0.095 × 98 = 9.31g/dm3